OpenStudy (anonymous):

my last question helppp?! Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (accidentalaichan):

x = number of votes Maya received: this is what you want to solve for y = total number of votes cast y is given to you = 125,000 Solve for x by plugging in y: x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 x = 93,750 = Number of votes Maya received

OpenStudy (anonymous):

your defff the best

OpenStudy (accidentalaichan):

Not the best, just in the same class ^^ Who's your Algebra teacher?

OpenStudy (anonymous):

mrs. bailey i think

OpenStudy (accidentalaichan):

Brittni Bailey? My teacher is Mrs Stefani :)

OpenStudy (anonymous):

did you take the practice test for module 2 ?

OpenStudy (anonymous):

& i think so ;)

OpenStudy (accidentalaichan):

Yes I have & Ms Bailey happens to be my choir teacher. I think she said something about teaching Algebra...

OpenStudy (anonymous):

help me ?! & really ? such a small world :0

OpenStudy (accidentalaichan):

I'll help you :3

OpenStudy (anonymous):

yaaay okay uhm do i just give you the question here ?

OpenStudy (accidentalaichan):

you can message me